Raner, Harris & Chan is a consulting firm that specializes in information systems for medical and dental clinics. The firm has two offices-one in Chicago and one in Minneapolis. The firm classifies the direct costs of consulting jobs as variable costs. A contribution format segmented income statement for the companys most recent year is given: Assume that Minneapolis sales by major market are: Market Minneapolin Medical Dental Sales Variable expenses Contribution margin Traceable fixed expenses Market segment margin Common fixed expenses not traceable to markets office segment margin $600,000 1008 $400,000 1008 $200,000 100 360,000 6 260.000 658100,00050 240,000 408 140,000 35 100,000 s0 72,000 124 20,000 59 52,000 26. 168,000 28 $120,000 30 48,000 24 18,000 3 $150,000 25 The company would like to initiate an intensive advertising campaign in one of the two market segments The campaign would cost $8,000. Marketing studies indicate that such a campaign would increase sales in the Medical market by $70,000 or increase sales in the Dental market by $60,000 during the next month. Required 1. How much would the companys profits increase (decreose) if it implemented the advertising campaign in the Medical Market? 2. How much would the companys profits increase (decrease) if it implemented the advertising campaign in the Dental Market? 3. In which of the markets would you recommend that the company focus its advertising campaign?
Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 How much would the companys profits increase (decrease) if it implemented the advertising campaign in the Medical Market? profits Required 2>
Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 How much would the companys profits increase (decrease) if it implemented the advertising campaign in the Dental Market? Companys profts C Required 1 Required 3
Answer #1

The points to be noted is VCR(variable cost ratio) and the net increase in profit

CASE 1: The advertisement is initiated in the medical market

VCR of the medical market = 65%

the increase in sales will be 70000

sales = $470 000 (400 000 + 70 000)

variable cost (65%) = $305 500

contribution (35% = $164 500

fixed expens = $28 000

profit   = $ 136 500

the difference made by advertisement = 136500 - 120000 =16500

CASE 2; the advertisement was initiated in the dental market

VCR of the dental market = 50%

sales = $260 000

VCR(50%) = $130 000

contribution = $130 000

fixed expenses = $60 000

profit = 70000

so the difference made by the advertisement =(70000) - 48000 = 22000

since profit is more in the dental market the suggestable way is to proceed investing in the dental market

note;

  • advertisement expenses are considered to be in the fixed expenses
  • there was more profit in the dental area despite having lower sales is because the VCR was lower in the dental market
  • this can be also calculated by just taking sales as 70000/60000 and subtracting VCR from them and deducting the advertisement expenses since the company has got past the break-even point